A plane flies 1500 miles against the wind in 3 hours and 45 minutes. the return trip with the wind takes 3 hours. What is the speed of the wind and the speed of the airplane?
I need a system of equations in two variables.

Answer by scott8148(6628)

p - w = 1500 / 3.75

p + w = 1500 / 3